log4(x-1)+log2(3x+4)

is it a trick because it is not equal to anything?
also we cannot use calculators
thanks,GodBless

Answer by Jk22(389)

log4(x-1)+log2(3x+4) = log4(x-1)+log(3x+4)/log(2)

= log4(x-1)+log4(3x+4)*log(4)/log(2)

= log4(x-1)+2log4(3x+4)

= log4((x-1)(3x+4)^2)